# On the torsion growth in quadratic number fields for elliptic curves defined over the rationals

### v1

@@ 5.2. Ordinary Reduction with Trivial Wild Inertia Action
$\ell=7$. Again, [14] studies this case and shows that $\overline{\rho}_{E,7}(G_{\mathbb{Q}})$ is conjugate in $\mathop{\mathrm{GL}}\nolimits_{2}(\mathbb{F}_{7})$ to a group from a list of 16 possible groups (see [14, Theorem 1.5]). Sutherland studies these groups and determines in each case the index of the largest subgroup that fixes a nonzero vector in $\mathbb{F}_{7}^{2}$; this quantity coincides with the degree of the minimal extension $K/\mathbb{Q}$ such that $E$ has a rational point of $7$-torsion.
@@
In [12, Table 3, p.65] we can find the list of these 16 groups along with these indices. There are only two of them where this degree is $2$, specifically those labeled as 7B.1.6, 7B.6.1. Let us examine each of these cases:
@@
    $\left\{\begin{pmatrix}6&0\\ 0&4\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\}$
@@
    $\left\{\begin{pmatrix}6&0\\ 0&6\end{pmatrix},\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\}$
@@
    $\left\{\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&6\end{pmatrix},\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\begin{pmatrix}1&0\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    $\left\{\begin{pmatrix}1&1\\ 0&3\end{pmatrix},\begin{pmatrix}1&4\\ 0&2\end{pmatrix},\begin{pmatrix}1&6\\ 0&6\end{pmatrix},\begin{pmatrix}1&5\\ 0&4\end{pmatrix},\begin{pmatrix}1&2\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    $\left\{\begin{pmatrix}1&3\\ 0&3\end{pmatrix},\begin{pmatrix}1&5\\ 0&2\end{pmatrix},\begin{pmatrix}1&4\\ 0&6\end{pmatrix},\begin{pmatrix}1&1\\ 0&4\end{pmatrix},\begin{pmatrix}1&6\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    $\left\{\begin{pmatrix}1&2\\ 0&3\end{pmatrix},\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\begin{pmatrix}1&5\\ 0&6\end{pmatrix},\begin{pmatrix}1&3\\ 0&4\end{pmatrix},\begin{pmatrix}1&4\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    $\left\{\begin{pmatrix}1&4\\ 0&3\end{pmatrix},\begin{pmatrix}1&2\\ 0&2\end{pmatrix},\begin{pmatrix}1&3\\ 0&6\end{pmatrix},\begin{pmatrix}1&6\\ 0&4\end{pmatrix},\begin{pmatrix}1&1\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    $\left\{\begin{pmatrix}1&5\\ 0&3\end{pmatrix},\begin{pmatrix}1&6\\ 0&2\end{pmatrix},\begin{pmatrix}1&2\\ 0&6\end{pmatrix},\begin{pmatrix}1&4\\ 0&4\end{pmatrix},\begin{pmatrix}1&3\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
    $\left\{\begin{pmatrix}1&6\\ 0&3\end{pmatrix},\begin{pmatrix}1&3\\ 0&2\end{pmatrix},\begin{pmatrix}1&1\\ 0&6\end{pmatrix},\begin{pmatrix}1&2\\ 0&4\end{pmatrix},\begin{pmatrix}1&5\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$
@@ 5.3. Supersingular Reduction
In this case, the image by $\overline{\rho}_{E,\ell}$ of the inertia group $I_{\ell}$ is an non-split Cartan subgroup (cf. [10, Section 1.9]). Consider a matrix
@@
If $a=1$ and $b\neq 0$, then $b\varepsilon y=0$ and $bx=0$, which implies $x=y=0$, contradicting the assumption that $(x,y)^{t}\not=(0,0)^{t}$. If $a\neq 1$, then $y=-bx/(a-1)$, and substituting into the other equation, we have
@@ 6. Conclusion and final remarks
We have then proved our original conjecture.
@@
While exploring the distinctive case $\ell=3$ we have also been able to prove some interesting remarks which give us a fuller picture of the phenomenon.
Proposition 10 (See Proposition 8).
Let $E/\mathbb{Q}$ be an elliptic curve and $K:=\mathbb{Q}(\sqrt{d})/\mathbb{Q}$ be a quadratic extension such that there exists a point $P\in E[3]$ satisfying that $P\in E(K)\setminus E(\mathbb{Q})$. Let us assume $3|d$ and $E$ has good reduction at 3, then
$E_{\mathrm{tors}}(K)=E_{\mathrm{tors}}(\mathbb{Q})\times\mathcal{C}_{3},$
where $\mathcal{C}_{3}$ denotes the cyclic group with $3$ elements.
@@ 7. Appendix: Matrix groups appearing in Section 6
    $\left\{\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&4\end{pmatrix},\;\begin{pmatrix}1&2\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&2\\ 0&4\end{pmatrix},\;\begin{pmatrix}1&3\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&3\\ 0&4\end{pmatrix},\right.$
@@
    $\left.\begin{pmatrix}4&3\\ 0&2\end{pmatrix},\ \begin{pmatrix}4&3\\ 0&3\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&2\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&3\end{pmatrix}\right\}$
@@
    $\left\{\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&3\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&4\end{pmatrix},\right.$
@@
    $\left.\begin{pmatrix}4&3\\ 0&1\end{pmatrix},\;\begin{pmatrix}4&3\\ 0&2\end{pmatrix},\;\begin{pmatrix}4&3\\ 0&3\end{pmatrix},\;\begin{pmatrix}4&3\\ 0&4\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&1\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&2\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&3\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&4\end{pmatrix}\right\}$
@@
    $\left\{\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&4\end{pmatrix},\;\begin{pmatrix}1&2\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&2\\ 0&2\end{pmatrix},\right.$
@@
    $\left.\begin{pmatrix}6&6\\ 0&2\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&4\end{pmatrix}\right\}$
@@
    $\left\{\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&5\end{pmatrix},\;\begin{pmatrix}1&0\\ 0&6\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\;\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\right.$
@@
    $\left.\begin{pmatrix}6&6\\ 0&3\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&4\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&5\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&6\end{pmatrix}\right\}$

### v2

@@ 5.2. Ordinary good reduction or multiplicative reduction, with trivial wild inertia action
Case 5.2.II: $\ell=7$. Again, [17] studies this case and shows that $\overline{\rho}_{E,7}(G_{\mathbb{Q}})$ is conjugate in $\mathop{\mathrm{GL}}\nolimits_{2}(\mathbb{F}_{7})$ to a group from a list of 16 possible groups (see [17, Theorem 1.5]). Sutherland studies these groups and determines in each case the index of the largest subgroup that fixes a nonzero vector in $\mathbb{F}_{7}^{2}$; this quantity coincides with the degree of the minimal extension $K/\mathbb{Q}$ such that $E$ has a rational point of $7$-torsion.
@@
In [15, Table 3, p.65] we can find the list of these 16 groups along with these indices. There are only two of them where this degree is $2$, specifically those labeled as 7B.1.6, 7B.6.1. Let us examine each of these cases:
@@
    $\left\{\begin{pmatrix}6&0\\ 0&4\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\}$.
@@
    $\left\{\begin{pmatrix}6&0\\ 0&6\end{pmatrix},\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right\}$.
@@
    $\left\{\begin{pmatrix}1&0\\ 0&3\end{pmatrix},\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\begin{pmatrix}1&0\\ 0&6\end{pmatrix},\begin{pmatrix}1&0\\ 0&4\end{pmatrix},\begin{pmatrix}1&0\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$,
    $\left\{\begin{pmatrix}1&1\\ 0&3\end{pmatrix},\begin{pmatrix}1&4\\ 0&2\end{pmatrix},\begin{pmatrix}1&6\\ 0&6\end{pmatrix},\begin{pmatrix}1&5\\ 0&4\end{pmatrix},\begin{pmatrix}1&2\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$,
    $\left\{\begin{pmatrix}1&3\\ 0&3\end{pmatrix},\begin{pmatrix}1&5\\ 0&2\end{pmatrix},\begin{pmatrix}1&4\\ 0&6\end{pmatrix},\begin{pmatrix}1&1\\ 0&4\end{pmatrix},\begin{pmatrix}1&6\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$,
    $\left\{\begin{pmatrix}1&2\\ 0&3\end{pmatrix},\begin{pmatrix}1&1\\ 0&2\end{pmatrix},\begin{pmatrix}1&5\\ 0&6\end{pmatrix},\begin{pmatrix}1&3\\ 0&4\end{pmatrix},\begin{pmatrix}1&4\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$,
    $\left\{\begin{pmatrix}1&4\\ 0&3\end{pmatrix},\begin{pmatrix}1&2\\ 0&2\end{pmatrix},\begin{pmatrix}1&3\\ 0&6\end{pmatrix},\begin{pmatrix}1&6\\ 0&4\end{pmatrix},\begin{pmatrix}1&1\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$,
    $\left\{\begin{pmatrix}1&5\\ 0&3\end{pmatrix},\begin{pmatrix}1&6\\ 0&2\end{pmatrix},\begin{pmatrix}1&2\\ 0&6\end{pmatrix},\begin{pmatrix}1&4\\ 0&4\end{pmatrix},\begin{pmatrix}1&3\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$,
    $\left\{\begin{pmatrix}1&6\\ 0&3\end{pmatrix},\begin{pmatrix}1&3\\ 0&2\end{pmatrix},\begin{pmatrix}1&1\\ 0&6\end{pmatrix},\begin{pmatrix}1&2\\ 0&4\end{pmatrix},\begin{pmatrix}1&5\\ 0&5\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right\}$.
@@ 5.3. Supersingular reduction
In this case, the image by $\overline{\rho}_{E,\ell}$ of the inertia group $I_{\ell}$ is an non-split Cartan subgroup (cf. [13, Section 1.9]). Consider a matrix
@@
If $a=1$ and $b\neq 0$, then $b\varepsilon y=0$ and $bx=0$, which implies $x=y=0$, contradicting the assumption that $(x,y)^{t}\not=(0,0)^{t}$. If $a\neq 1$, then y=bx/(a1)y=-bx/(a-1), and substituting into the other equation, we have
@@
5.4. Main result for $\ell=5,7$.
This last argument finishes the proof of Theorem 3. However, our techniques allow us to prove a stronger version of this result, as stated in the introduction (and including some additional information on $\ell=3$).
Theorem 4. Let $E/\mathbb{Q}$ be an elliptic curve. Assume $K$ is a quadratic number field such that there exists $P\in E(K)[\ell]\setminus E(\mathbb{Q})$, with $\ell\geq 3$ prime.
  1. (1)
    For every prime $p\neq\ell$ that ramifies in $K$, $E$ has additive reduction at $p$, i.e. $p^{2}|N_{E}$.
  2. (2)
    If $p=\ell>3$ ramifies in $K$, $E$ has additive reduction at $p$, i.e. $p^{2}|N_{E}$.
Proof.
By hypothesis,
$\mathbb{Q}\subsetneq K\cap\mathbb{Q}(E[\ell])\subset K\quad\text{and}\quad[K:\mathbb{Q}]=2.$
The existence of $P$ implies that $K\cap\mathbb{Q}(E[\ell])\not$$\mathbb{Q}$, thus $K\subset\mathbb{Q}(E[\ell])$, which implies that $p$ ramifies in $\mathbb{Q}(E[\ell])$.
Let us prove the first part of the theorem. We have that, by hypothesis, $p\not$$\ell$. From the Néron–Ogg–Shafarevich criterion (see [14, VII, Theorem 7.1]), we can conclude that $E$ has bad reduction at $p$. By way of contradiction, suppose that $E$ has multiplicative reduction at $p$.
Let $\mathfrak{P}|p$ be a prime above $p$ in $\mathbb{Q}(E[\ell])$. Let $I(\mathfrak{P}|p)$ denote the inertia group associated with this prime, and let
$L=\mathbb{Q}(E[\ell])^{I(\mathfrak{P}|p)}$
be the maximal unramified subextension at $p$ of $\mathbb{Q}(E[l])|\mathbb{Q}$. Since $p$ ramifies in $K$, we have the following diagram:
$\mathbb{Q}(E[\ell])$$L{L}K{K}{\mathbb{Q}}$$\scriptstyle{2}$$\scriptstyle{I(\mathfrak{P}|p)}$$\scriptstyle{2}$
By Proposition 10, $I(\mathfrak{P}|p)$ has order $\ell\geq 3$. Since $\ell>2$, the diagram yields a contradiction.
The second part of the theorem was already proven in this section. ∎
@@ 6. Conclusion and final remarks
Merging together Proposition 2, Theorem 2 (proven in Sections 2 and 3), and Theorem 3 (proven in Section 5), we can state the following result:
@@
For $p>3$ (that is, Theorem 3), this theorem has already been proved with different techniques and in a more general context by Mentzelos Melistas (see [10, Theorem 1.5]).
Also, when $p>2$, the condition that $E$ has a $p$-torsion point in $E(K)\setminus E(\mathbb{Q})$, where $K=\mathbb{Q}(\sqrt{d})$, is equivalent to the condition that the quadratic twist $E_{d}$ has a rational $p$-torsion point. As one of the referees pointed out, one could try to prove Theorem 1 for $p>3$ relying on this fact. And, indeed, by makes use of some results of Olson [12], one can find an alternative proof in this case.
However, for $p>3$, the techniques used in the alternative proof given in Section 5 allowed us to prove Theorem 4, which is a stronger result.
While exploring the distinctive case $\ell=3$ we have also been able to offer some interesting remarks which give us a fuller picture of the phenomenon. Mainly:
Proposition 8. Let $E/\mathbb{Q}$ be an elliptic curve and $K=\mathbb{Q}(\sqrt{d})$ be a quadratic extension such that there exists a point $P\in E[3]$ satisfying that $P\in E(K)\setminus E(\mathbb{Q})$. Let us assume $3|d$ and $E$ has good reduction at 3, then
$E_{\mathrm{tors}}(K)=E_{\mathrm{tors}}(\mathbb{Q})\times\mathcal{C}_{3}.$
@@ 7. Appendix: Matrix groups appearing in Section 5
[…]
@@
    $\left.\begin{pmatrix}4&3\\ 0&2\end{pmatrix},\ \begin{pmatrix}4&3\\ 0&3\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&2\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&3\end{pmatrix}\right\}$.
@@
[…]
@@
    $\left.\begin{pmatrix}4&3\\ 0&1\end{pmatrix},\;\begin{pmatrix}4&3\\ 0&2\end{pmatrix},\;\begin{pmatrix}4&3\\ 0&3\end{pmatrix},\;\begin{pmatrix}4&3\\ 0&4\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&1\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&2\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&3\end{pmatrix},\;\begin{pmatrix}4&4\\ 0&4\end{pmatrix}\right\}$.
@@
[…]
@@
    $\left.\begin{pmatrix}6&6\\ 0&2\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&4\end{pmatrix}\right\}$.
@@
[…]
@@
    $\left.\begin{pmatrix}6&6\\ 0&3\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&4\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&5\end{pmatrix},\;\begin{pmatrix}6&6\\ 0&6\end{pmatrix}\right\}$.
Conflict of Interest: Not Applicable.